OpenStudy (anonymous):

Select the correct symbol to complete the given statement, 4/5 times 20 is < > = to 20 5/4 times 20 is < > = to 20 5/5 times 20 is< > = to 20

OpenStudy (anonymous):

plz reply

OpenStudy (er.mohd.amir):

4/5*20=16 5/4*20=25 5/5*20=20

OpenStudy (anonymous):

no i have to know is it < > or =

OpenStudy (harley1312):

Which one do you think it is?

OpenStudy (anonymous):

help plz

OpenStudy (er.mohd.amir):

16 is > or < to 20 > greater < lester symbol

OpenStudy (er.mohd.amir):

1. < 2. > 3.=